A particle of mass m moves according to laws x=A·cosωt and y=B·sinωt. Find the force acting on thisparticle in each point of its trajectory.


1
Expert's answer
2021-06-16T15:38:32-0400
"\bold{r}=a \cos ωt\bold{i }+ b \sin ωt\bold{j}\\
\bold{v}=b\omega \cos ωt\bold{j }-a\omega \sin ωt\bold{i}\\
\bold{a}=-\omega^2(a \cos ωt\bold{i }+ b \sin ωt\bold{j})=-\omega^2\bold{r}\\
\bold{F}=-m\omega^2(a \cos ωt\bold{i }+ b \sin ωt\bold{j})=-m\omega^2\bold{r}\\"